### v1

# AI Chatbots as Professional Service Agents: Developing a Professional Identity

Abstract
With the rapid expansion of large language model (LLM) applications, there is an emerging shift in the role of LLM-based AI chatbots from serving merely as general inquiry tools to acting as professional service agents. However, current studies often overlook a critical aspect of professional service agents: the act of communicating in a manner consistent with their professional identities. This is of particular importance in the healthcare sector, where effective communication with patients is essential for achieving professional goals, such as promoting patient well-being by encouraging healthy behaviors. To bridge this gap, we propose LAPI (LLM-based Agent with a Professional Identity), a novel framework for designing professional service agent tailored for medical question-and-answer (Q&A) services, ensuring alignment with a specific professional identity. Our method includes a theory-guided task planning process that decomposes complex professional tasks into manageable subtasks aligned with professional objectives and a pragmatic entropy method designed to generate professional and ethical responses with low uncertainty. Experiments on various LLMs show that the proposed approach outperforms baseline methods, including few-shot prompting, chain-of-thought prompting, across key metrics such as fluency, naturalness, empathy, patient-centricity, and ROUGE-L scores. Additionally, the ablation study underscores the contribution of each component to the overall effectiveness of the approach.
1 Introduction
AI chatbots have gained great popularity worldwide to serve as professional assistants and provide services across various sectors Dam et al. (2024); Xi et al. (2023). Integrating AI chatbots is a cost-effective strategy for organizations, delivering value to customers while reducing operational expenses. For instance, automation in customer interactions is projected to save retailers approximately \$439 billion annually by 2023, as they replace many human-operated customer service roles with chatbots Williams (2019). The emergent of generative AI, particularly large language models (LLMs), is revolutionizing the functionality of chatbots by significantly enhancing their interactive communication capabilities. LLMs, such as OpenAI’s GPT-series and Llama, enabling chatbots to understand and generate natural language with remarkable fluency and context-awareness  Achiam et al. (2023); Touvron et al. (2023a).
Recently, there is a trend of a significant transition in chatbot applications, shifting from general inquiry tools to professional service agents capable of delivering specialized support. The increasingly prevalent human-AI interactions in high-stake industries (e.g., healthcare and finance) require professional service agents to offer a more professional and functional experience than simple knowledge-enhanced inquiry tools Safi et al. (2020). This leads to a critical challenge in human-AI communication that has not been well addressed in existing studies, namely how AI chatbots can communicate with humans and answer questions in a manner consistent with the chatbots’ professional identity. Beyond simply responding to user queries, professional service agents always have a defined professional identity, such as that of a virtual doctor, which requires their responses to be aligned with professional goals, such as supporting patients’ health behavior change (e.g., treatment or medication adherence and reducing substance abuse)Clark and Bailey (2024). However, existing AI chatbots primarily emphasize information retrieval and question answering, overlooking the professional intention and goals that are integral to professional fields  Lee et al. (2024); Singhal et al. (2025).
To develop a professional identity, an AI chatbot should have the ability to think and act like a professional. Firstly, the ability to analyze and break down complex professional tasks into more manageable components is essential. In contrast to mathematical and logical problems, which typically have clear chains of thought and definitive answers Wei et al. (2022), many question-and-answer (Q&A) tasks in a service scenario are open-ended and require free-form responses Megahed et al. (2024); Singhal et al. (2025). The challenge lies in effectively guiding these chatbots to adopt a professional thought process. Secondly, given their role as frontline professionals, AI chatbots must have the competence to effectively address user inquiries and deliver reliable services. However, it should be noted that LLMs are non-deterministic, meaning that the same input can generate different outputs, some of which may be incorrect or hallucinated Farquhar et al. (2024). These drawbacks can result in occasional inaccurate answers or hallucination problems, which is fatal in the high-stakes field Ji et al. (2023). Interactions with such chatbots have the potential to trigger undesirable behaviors in users, such as noncompliance among patients, which can negatively impact both service providers and users.
Motivated by to urgent need to provide professional service to users through LLM-based chatbots, our goal is to propose a novel framework, i.e., LAPI (LLM-based Agent with a Professional Identity), which aims to enhance the effectiveness of LLM-based chatbots in serving as professional service agents, particularly in the context of medical Q&A services. To address the aforementioned challenges, we consider a subfield of linguistics, namely pragmatics, which how language is used for social interactions and the attainment of goals Ciccia and Turkstra (2002) and the use of information from context to determine the effective means of achieving a given goal Grice (1975). From a pragmatic perspective, AI chatbots should provide useful responses and remove meaningless or incorrect information. The proposed approach involves two key steps: first, the comprehension and decomposition of complex professional tasks into manageable subtasks through a theory-guided task planning process; and second, the iterative updating and eventual generation of professional and ethical responses using a pragmatic entropy method. Specifically, we propose a theory-guided task planning process that employs a conceptual framework in the healthcare domain—namely, the Health Belief Model (HBM)—to guide AI chatbots in how they think and act to support interventions aimed at encouraging healthy behaviors. Furthermore, we propose a pragmatic entropy method that generates professional and ethical responses, with the objective of maximizing the clarity, empathy, and relevance of the chatbot’s outputs while ensuring alignment with ethical standards and professional goals.
In general, our contributions are:
  • We introduce a novel framework for developing a professional identity for an LLM-based chatbot, with the objective of employing the chatbot as a professional service agent for medical Q&A services. Specifically, we propose a pragmatic entropy method for the generation of professional and ethical responses. Meanwhile, we provide an effective way to utilize a domain theory to guide task planning process.
  • Experiments on various real medical Q&A setting show that our method significantly improves response quality, providing more accurate, empathetic, and professional answers compared to baseline approaches.
2 Research Background
2.1 Medical AI Chatbots
The medical field represents a domain in which the implementation of AI chatbots is gaining traction as a means of facilitating access to information from the patient perspective and alleviating the workload on doctors Safi et al. (2020). A variety of methods and techniques have been employed in the development of medical chatbots for diverse objectives, including statistical methods, deep learning, and LLMs Mihailidis et al. (2008); Jothi et al. (2022); Lee et al. (2024). These bots have been developed for diverse applications, including various aspects of health and medicine. These range from chatbots that support general health and activities of daily living to chatbots that assist specific disease management, such as heart disease and cancerWong et al. (2011); Lee et al. (2024). Prior to the advent of LLMs, conventional chatbots exhibited constrained contextual and language understanding, resulting in inaccurate responses and suboptimal interaction in a human-like manner. Recently, significant advancements have been made in the development of LLM-based medical chatbots, which are designed to facilitate human-like interactions with patients Singhal et al. (2025).
2.2 Professional LLM-based Chatbots
Chatbots have stepped into the age of LLM empowerment. The rapid development of LLMs enhances the capabilities of chatbots by improving their conversational proficiency and facilitating human-like interactions. ChatGPT exemplifies this category of chatbot, optimized to generate natural, human-like dialogue. Other widely used LLMs include GPT4, Llama, Mistral, ChatGLM2, and Gemma Achiam et al. (2023); Touvron et al. (2023a); Jiang et al. (2023); Zeng et al. (2024); Mesnard et al. (2024). Chatbots have been employed in customer service to address customer inquiries regarding products or services, either as an initial point of contact or as an alternative to speaking with a human representative Hong et al. (2024). In addition to healthcare, LLM-based chatbots have gradually become powerful tools in areas such as retail, education, research, and many others Megahed et al. (2024); Benzinho et al. (2024); Giudici et al. (2024). To improve chatbot’s ability on professional tasks, such as programming education and home designing, various knowledge-enhanced strategies, such as fine-tuning, prompt engineering, and Retrieval-Augmented Generation (RAG), have been employed Lewis et al. (2020).
3 Theoretical Background of Professional Service Agents
3.1 Professional Service Agents
From a philosophical standpoint, the term ”agent” traditionally describes entities possessing desires, beliefs, intentions, and the capacity to action Xi et al. (2023); Davidson (1963). Human agents, for instance, are frequently employed by companies to handle customer service operations, focusing on addressing customer queries and ensuring satisfaction through effective communication and problem-solving skills. With the progression of AI, the concept of an ”agent” has expanded to describe AI entities that are characterized by intelligent behavior, such as autonomy, proactiveness, and social capabilities. AI agents, particularly chatbots powered by large language models (LLMs), have demonstrated remarkable versatility across various scenarios, attributable to their extensive inherent knowledge and human-like communication abilities Huang et al. (2024); Chong et al. (2021).
LLM-based agents are typically structured around three key components. Agent Core serves as the main brain of the system, typically utilizing an LLM with general-purpose capabilities. Task Planning is responsible for decomposing complex tasks into manageable subtasks, thereby enhancing the agent’s reasoning and reliability of responses. Actions of a chatbot are communication with humans. A professional AI agent’s primary actions are aimed at high-level information transmission to fulfill a robot’s intent, a process referred to as “intention action.” A typical example of such agents is applications in the healthcare domain, where empathy-aware agents assist in diagnostics and knowledge retrieval to provide a user-friendly service Mao et al. (2022).
Figure 1: Distinction Between AI Chatbots as Inquiry Tools and Professional Service Agents
Responding to the trend of chatbots shifting from general inquire tools to professional service agents, we identify the significant distinction between AI-chatbots as inquiry tools and as professional service agents, as shown in Figure 1. Professional identity is a crucial attribute of a professional service agent, significantly shaping the agent’s goals, task planning, and action execution. Chatbots with professional identity have distinct professional intentions and goals, which in turn influence their task planning and action execution. Using medical Q&A as an example, inquiry tools primarily focus on answering medical questions Safi et al. (2020). In contrast, professional service agents serve two purposes: responding to medical inquiries and promoting patient well-being by encouraging healthy behaviors, aligning with their professional identity (e.g., doctor). For task planning, the decomposition of complex professional tasks should adhere to established professional standards, informed by either theories or expert guidelines. Regarding actions, the chatbot’s responses should align with professional goals. This alignment ensures that the chatbot not only addresses patient questions but also offers suggestions and persuasive guidance to foster healthy behaviors.
Figure 2: Overview of LLM-based Agent with a Professional Identity (LAPI). The patient’s query is first assigned weights for the two HBM domains—Perceived Threat and Perceived Effectiveness of Health Behavior. An initial prompt is then generated based on these weights, followed by the addition of pragmatic rules. Multiple outputs are generated by the LLM, and pragmatic entropy optimization is applied to obtain the most appropriate response.
3.2 Health Belief Model
The Health Belief Model (HBM) is a widely used conceptual framework in health behavior research, which supports interventions aimed at encouraging healthy behaviorsChampion and Skinner (2008). Initially proposed in the 1950s, the model has evolved in response to practical public health concern and healthcare management. Rooted in a well-established body of psychological and behavioral theories, the HBM posits that health-related behavior depends primarily influenced by two factors: (1) the individual’s desire to avoid illness or recover from it if already affected; and (2) the belief that a particular health action care prevent or mitigate the illness Janz and Becker (1984).
The HBM comprises four key constructions: Perceived Susceptibility, Perceived Severity, Perceived Benefit, and Perceived Barrier. These constructs can be broadly categorized into two main aspects: Perceived Threat (including Perceived Susceptibility and Perceived Severity) and Perceived Effectiveness of Health Behavior (including Perceived Benefit and Perceived Barrier). Perceived Susceptibility refers to an individual’s belief about the likelihood of contracting a particular condition. Perceived Severity refers to the belief regarding the seriousness of contracting an illness or the consequences of leaving it untreated. Perceived Benefit relates to the belief in the effectiveness of the recommended action to mitigate the disease threat. Perceived Barrier refers to the belief regarding the potential negative aspects or costs associated with the recommended health action. Individuals tend to weigh the anticipated benefits of an action against the perceived barriers, resulting in a belief in the effectiveness of the health behavior. These four constructions interact to guide individuals toward a preferred course of action.
The HBM has been employed to inform health intervention designs aimed at promoting healthy behaviors Orji et al. (2012), particularly as patient-centric (proactive) healthcare gains popularity. Previous research has shown that interventions grounded in established theories and models tend to be more effective than those developed based on intuition Glanz (1997).
4 Methods
4.1 Problem Formulation
Given a patient’s question $q$ and a set of predefined pragmatic rules aligning with a professional identity $R=\{r_{1},r_{2},r_{3},\dots,r_{m}\}$, the objective is to generate a response $s$ that not only provides an professional answer to the query but also adheres to these established rules, thereby reflecting the professional identity (i.e., the doctor in our experiment setting). The desired response $s^{*}$ should satisfy the following condition: it adheres to the pragmatic rules with a minimal level of uncertainty in the response distribution. The problem can be formally expressed as:
$s^{*}=\text{LLM}(q,R,\tau,\epsilon)$ (1)
where $\tau$ is the satisfaction rate threshold, and $\epsilon$ is the pragmatic entropy threshold. The final response $s^{*}$ is selected when the satisfaction rate $P(C_{\text{all}}|S)\geq\tau$ and the pragmatic entropy $PE(S)\leq\epsilon$, ensuring both accuracy and alignment with the predefined professional characteristics.
Figure 2 illustrates the proposed framework, i.e., LLM-based Agent with a Professional Identity (LAPI), which includes a theory-guided task planning process and a pragmatic entropy method. In the following sections, we introduce details on our specific implementation of these two components.
4.2 Theory-guided Task Planning
We employ the Health Belief Model (HBM) as a guiding framework to structure our design, decomposing the professional goal of promoting healthy behavior into manageable sub-tasks. These sub-tasks include: fostering individuals’ accurate perception of disease risk, aiding individuals in developing a realistic understanding of the consequences associated with a condition or recommended action, enhancing individuals’ awareness of the potential benefits of adopting healthy behaviors, and offering strategies to overcome perceived barriers to implementing these behaviors.
The task planning process begins by assessing the relevance of the user’s query to the two core categories of HBM: Perceived Threat and Perceived Effectiveness of Health Behavior. To achieve this, we first prompt the LLM to estimate the proportions of the query related to these two categories.
Figure 3: Prompt used to assign weights to a question based on its relevance to the HBM categories.
Based on the prompt shown in Figure 3, the LLM assesses the query’s relevance to each category and assigns a corresponding weight to each. Next, we generate the initial prompt $p_{\text{initial}}$ for the LLM by constructing two specific prompts: one fully focused on the Perceived Threat category, and another fully focused on Perceived Effectiveness of Health Behavior. These prompts are pre-configured to ensure that the LLM generates an initial prompt that accurately reflects the correct placement of the question within the HBM theory framework. The content of these prompts is in Appendix A.
4.3 Pragmatic Entropy
Pragmatic entropy aims to assist the AI chatbot generate professional and ethical response by 1) iteratively optimizing a given prompt $p$ to generate responses that maximize the proportion of outputs satisfying all predefined pragmatic rules $R=\{r_{1},r_{2},r_{3},...,r_{m}\}$, and 2) selecting responses with minimized uncertainty from the category that meet all pragmatic rules. Pragmatic entropy measures the variability in the distribution of responses across different rule-satisfaction categories, with lower entropy indicating more focused outputs that align with the desired rules. By minimizing entropy, the method ensures stable and reliable generation, improving alignment with professional goals.
Response Evaluation.
Given a prompt $p$, the LLM generates a set of $n$ responses by sampling from its output distribution with temperature 1, which ensures a balance between randomness and determinism in the generated outputs:
$S=\text{LLM}(p,\text{temperature}=1)$ (2)
where $S=\{s_{1},s_{2},\dots,s_{n}\}$ represents the generated responses. The temperature parameter controls the randomness of the sampling process. A temperature of $1$ maintains the default probability distribution of the model, allowing diverse but coherent outputs.
Each response $s\in S$ is classified into categories $C_{j}$ based on its satisfaction of specific subsets of the pragmatic rules $R$. This classification process is judged by the Large Language Model (LLM) itself. The following metrics are computed to evaluate the quality of the prompt:
  1. 1.
    Satisfaction Rate: The proportion of responses satisfying all rules is given by:
    $P(C_{\text{all}}|S)=\frac{\left|\{s_{i}\in S\mid s_{i}\in C_{\text{all}}\}\right|}{n}$ (3)
    where |$\left|\{s_{i}\in S\mid s_{i}\in C_{\text{all}}\}\right|$ denotes the number of responses in $S$ that fall into category $C_{\text{all}}$, and $n$ is the total number of generated responses.
  2. 2.
    Pragmatic Entropy: The entropy of the output distribution across all categories $C_{j}$ is computed as:
    $PE(S)=-\sum_{C_{j}}P(C_{j}|S)\log P(C_{j}|S)$ (4)
    where $P(C_{j}|S)=\frac{\left|\{s_{i}\in S\mid s_{i}\in C_{j}\}\right|}{n}$ represents the proportion of responses in $S$ that fall into category $C_{j}$.
Optimization Process
To refine the prompt iteratively, the process consists of generating an improved prompt, evaluating it quantitatively, and iterating until either the desired criteria are met or the maximum number of iterations is reached.
To improve the prompt, the current prompt $p$ along with information on unmet conditions is fed into the large language model (LLM) to generate a refined version $p_{\text{new}}$:
$p_{\text{new}}=\text{LLM}(p,unmet\_conditions,\text{temperature}=0.3)$ (5)
where $p_{\text{new}}$ is the updated prompt. Unmet conditions refer to the missing pragmatic rules identified within the category that has the largest proportion in categories that have not satisfied all the rules. The parameter $\text{temperature}=0.3$ Moslem et al. (2023); Wassie (2024) is chosen to balance stability and flexibility in the prompt generation process.
The refined prompt $p_{\text{new}}$ is evaluated quantitatively by generating a new set of $n$ responses:
$S_{\text{new}}=\text{LLM}(p_{\text{new}},\text{temperature}=1).$ (6)
Each response $s\in S_{\text{new}}$ is classified into categories $C_{j}$, and the satisfaction rate $P(C_{\text{all}}|S_{\text{new}})$ and pragmatic entropy $PE(S_{\text{new}})$ are recomputed.
If the satisfaction rate $P(C_{\text{all}}|S)$ meets or exceeds a predefined threshold $\tau$, and the pragmatic entropy $PE(S)$ is below a threshold $\epsilon$, the prompt $p$ is considered optimal, and the process terminates with $p^{*}=p$. Otherwise, the model proceeds to refine the prompt.
Algorithm 1 Pragmatic entropy optimization process
  Input: Current prompt $p$, pragmatic rules $R$, satisfaction threshold $\tau$, entropy threshold $\epsilon$, maximum iterations $T$.
  Output: Final response answer.
1:  Set iteration counter $t\leftarrow 0$;
2:  while $t<T$ do
3:     Generate response set $S$ using Eq.(2);
4:     Compute satisfaction rate $P(C_{\text{all}}|S)$ using Eq.(3);
5:     Compute pragmatic entropy $PE(S)$ based on Eq.(4);
6:     Select $s^{*}$ randomly from $S$;
7:     if $P(C_{\text{all}}|S)\geq\tau$ and $PE(S)\leq\epsilon$ then
8:        Return $s^{*}$;
9:     else
10:        Identify $unmet\_conditions$ from categories that have not satisfied all the rules;
11:        Get the updated prompt $p_{\text{new}}$ using Eq.(5);
12:        Set $p\leftarrow p_{\text{new}}$;
13:     end if
14:     Increment iteration counter $t\leftarrow t+1$
15:  end while
16:  Return $s^{*}$
5 Experiments
5.1 Experimental Settings
Models Approach Flu. Coh. Nat. Emp. Pat-C. Rec. Pre. Llama-2-7B-Chat Zero-Shot 3.90 3.60 2.95 2.10 2.45 0.39 0.07 Zero-Shot + Pragmatic rules 4.00 3.50 3.40 2.70 3.00 0.33 0.10 RAG-based Few-Shot 3.60 2.50 2.65 2.20 2.45 0.38 0.07 Chain-of-Thought 4.00 3.60 2.80 2.20 2.40 0.40 0.07 LAPI 4.00 3.60 3.05 2.90 3.10 0.35 0.07 Llama-2-13B-Chat Zero-Shot 3.75 3.60 2.85 2.25 2.65 0.38 0.08 Zero-Shot + Pragmatic rules 4.00 3.75 3.10 3.05 3.25 0.33 0.12 RAG-Based Few-Shot 3.60 2.65 2.75 2.35 2.25 0.31 0.07 Chain-of-Thought 3.65 3.65 2.95 2.10 2.50 0.39 0.07 LAPI 4.00 3.75 3.15 3.40 3.45 0.36 0.07 Llama-2-70B-Chat Zero-Shot 3.90 3.45 3.15 2.15 2.80 0.35 0.09 Zero-Shot + Pragmatic rules 4.00 3.75 3.40 2.80 3.35 0.35 0.08 RAG-Based Few-Shot 3.80 3.20 2.95 2.35 2.95 0.31 0.11 Chain-of-Thought 3.95 3.45 3.05 2.40 2.70 0.38 0.07 LAPI 4.00 3.75 3.50 3.35 3.35 0.36 0.07 GPT-3.5-Turbo Zero-Shot 4.00 3.60 3.30 2.35 3.15 0.25 0.15 Zero-Shot + Pragmatic rules 4.00 3.60 3.45 3.15 3.55 0.26 0.16 RAG-Based Few-Shot 4.00 3.45 3.25 2.10 3.30 0.26 0.24 Chain-of-Thought 4.00 3.70 3.10 2.30 3.10 0.29 0.16 Semantic Entropy 4.00 3.40 3.10 2.30 3.15 0.21 0.21 LAPI 4.00 3.85 3.45 3.50 3.55 0.30 0.13 GPT-4 Zero-Shot 4.00 3.70 3.05 2.10 2.90 0.35 0.08 Zero-Shot + Pragmatic rules 4.00 3.90 3.50 3.55 3.80 0.35 0.08 RAG-Based Few-Shot 4.00 3.65 3.25 2.30 3.25 0.36 0.11 Chain-of-Thought 4.00 3.70 3.10 2.35 2.85 0.37 0.08 Semantic Entropy 4.00 3.65 3.05 2.40 2.80 0.35 0.07 LAPI 4.00 3.90 3.95 3.85 3.80 0.38 0.07
Table 1: Evaluation results. The table presents the performance of different prompting approaches across various evaluation metrics. Flu., Coh., Nat., Emp., and Pat-C. represent Fluency, Coherence, Naturalness, Empathy, and Patient-Centricity, respectively. Rec. and Pre. refer to ROUGE-L Recall and Precision. Higher scores for the first six metrics (Flu., Coh., Nat., Emp., Pat-C., and Rec.) indicate better performance, while a lower score for Pre. reflects a more desirable outcome. The LAPI approach is often highlighted for its consistency and superior performance across most models and metrics. Bold: Best-performing paradigm per metric.
Dataset and Metrics
The dataset used in our experiments was collected from a reputable ophthalmology hospital and comprises 400 real-world patient inquiries paired with concise and professional answers provided by expert ophthalmologists. To support Retrieval-Augmented Generation (RAG), we utilized 300 patient-question-answer pairs as the knowledge base for retrieval. To assess the efficacy of the proposed method, a test set of 10 patient inquiries was randomly selected. This test set was used to evaluate the performance of the method in handling real-world medical question-and-answer scenarios.
To comprehensively evaluate the performance of our proposed method, we utilized two categories of metrics: LLM-based and non-LLM-based metrics. For LLM-based evaluation, we employed G-EVAL Liu et al. (2023), a state-of-the-art metric designed to assess various qualitative aspects of large language model outputs, including Fluency, Coherence, Naturalness, Empathy, and Patient-Centricity, where higher scores indicate better performance. Detailed descriptions of these five evaluation metrics can be found in Appendix B. G-EVAL has been shown to exhibit strong consistency with human evaluations Zhu et al. (2025); Gao et al. (2024), making it a reliable tool for qualitative assessment. For non-LLM-based evaluation, we adopted ROUGE-L recall and precision Lin (2004) as objective n-gram matching-based metrics. ROUGE-L Recall measures the extent to which the generated responses cover the content of the doctor-provided answers, with higher values indicating better coverage, while ROUGE-L Precision evaluates the informativeness of the generated responses, with higher values indicating that the generated responses contain more information.
Figure 4: Overview of GPT-3.5-Turbo’s responses using different methods to answer the patient query: ”After undergoing laser surgery for high myopia, will the degree of myopia increase again?” The text in green provides the direct answer to the question, the text in yellow encourages regular contact with the patient’s doctor and the text in blue offers actionable advice for the patient.
Baselines
We categorize the baseline methods into two groups: prompt design-based methods and quality inspection-based methods. For the prompt design-based methods, we explored strategies to guide the model’s reasoning and response generation. Specifically, we include the following: Zero-Shot, where the model generates responses directly without additional guidance; Zero-Shot + Pragmatic Rules, which enhances the zero-shot approach by incorporating explicit pragmatic rules tailored to professional contexts; RAG-based Few-Shot Ori et al. (2023); Guoxin et al. (2023), a retrieval-augmented generation strategy that retrieves relevant examples from a knowledge base to assist response generation; and Chain-of-Thought (CoT) Kojima et al. (2023), which encourages the model to generate step-by-step reasoning paths to improve response accuracy and relevance. For the quality inspection-based methods, we employed Semantic Entropy Farquhar et al. (2024), a technique designed to evaluate the confidence of the model regarding specific knowledge points in its generated responses. The confidence scores are then fed back into the model, prompting it to revise and improve its outputs.
Experiment Details
We evaluate our proposed method on five large language models (LLMs): Llama-2-7B-Chat, Llama-2-13B-Chat, Llama-2-70B-Chat Touvron et al. (2023b), GPT-3.5-Turbo (gpt-3.5-turbo-0125), and GPT-4 (gpt-4-0125-preview), selected for their diverse scales and capabilities, representing both open-source and closed-source models. To enhance the alignment of responses with professional and human-like communication, we introduced pragmatic rules used in both Zero-Shot+Pragmatic Rules and LAPI approaches, as detailed in Appendix C.
For the RAG-based Few-Shot method, we employed a retrieval-augmented generation strategy using a BERT encoder to retrieve the three most relevant question-answer pairs based on cosine similarity, which were then used as prompts to guide response generation. In the HBM + Pragmatic Entropy method, we set $n$ to 5, $\tau$ to 0.5, and $\epsilon$ to 0.8. For the Semantic Entropy method, due to the high text comprehension capability required for this method, Semantic Entropy was only applied to GPT-3.5-Turbo and GPT-4.
5.2 Experimental Results
Our proposed approach LAPI demonstrates superior performance across various models, as evidenced by the results in Table 1. From the results perspective, LAPI consistently outperforms other methods in terms of Fluency, Coherence, Naturalness, Empathy, Patient-Centricity, ROUGE-L Recall, and Precision. For instance, in the same model family (e.g., Llama-2-7B-Chat), LAPI provides the highest scores across multiple metrics highlighting the effectiveness of LAPI in generating more precise, empathetic, and coherent responses, tailored to professional service agent tasks. From the model perspective, we observe that LAPI performs even better on larger models such as GPT-3.5-Turbo and GPT-4. The enhanced textual comprehension capabilities of these models enable them to generate more accurate prompts and subsequently produce better results. LAPI benefits from this ability, yielding more contextually rich and reliable outputs.
The combination of pragmatic rules and the HBM framework allows LAPI to produce responses that are not only accurate and fluent but also deeply empathetic and patient-centric. By integrating step-by-step reasoning and continuous feedback mechanisms, our method ensures the delivery of high-quality, professional, and reliable responses. These results demonstrate the potential of LAPI to improve interactions in professional service applications, especially in the healthcare sector where empathy and clarity are paramount.
5.3 Ablation Study
We conducted an ablation study using GPT-3.5-Turbo as the base model to evaluate the effectiveness of each component in our proposed method. Table 2 presents the results of the ablation study, where we compare three different setups: the full LAPI method, the model without the HBM framework w/o HBM, and the model without the Pragmatic Entropy component w/o PE. As shown in the table, the full LAPI method outperforms both variants in all evaluation metrics, including Fluency, Coherence, Naturalness, Empathy, Patient-Centricity, ROUGE-L Recall, and Precision.
The results indicate that each component contributes positively to the overall performance. Specifically, removing the HBM framework (w/o HBM) leads to a slight decrease in the Coherence and Empathy metrics, while the absence of the Pragmatic Entropy (w/o PE) reduces the overall Naturalness and Empathy. These findings suggest that combining both the HBM framework and Pragmatic Entropy is more effective than using them separately, as the combined approach leverages the strengths of both components, resulting in more fluent, coherent, and empathetic responses tailored to professional service applications.
Approach Flu. Coh. Nat. Emp. Pat-C. Rec. Pre.
LAPI 4.00 3.85 3.60 3.50 3.70 0.29 0.13
w/o HBM 4.00 3.70 3.75 3.30 3.60 0.27 0.15
w/o PE 3.90 3.70 3.50 2.95 3.60 0.26 0.16
Table 2: Ablation study on GPT-3.5-turbo.
5.4 Case Study
To further illustrate the capabilities of our proposed method, we present a case study using GPT-3.5-Turbo to answer the patient query: ”After undergoing laser surgery for high myopia, will the degree of myopia increase again?”. Figure 4 provides a comparison of the responses generated by four different prompting strategies: LAPI, Chain-of-Thought, Zero-Shot, and RAG-Based Few-Shot. In the responses, the text in green represents the direct answer to the question, the text in yellow encourages regular contact with the patient’s doctor, and the text in blue offers actionable advice for the patient.
As observed in the case study, LAPI (left) outperforms methods like Chain-of-Thought, Zero-Shot, and RAG-Based Few-Shot. LAPI gives a more comprehensive and detailed response, directly answering the question, stressing the importance of regular ophthalmologist follow-ups, and offering advice on healthy eye habits. This showcases the benefit of integrating the HBM framework and Pragmatic Entropy. In contrast, other methods provide direct answers and some patient advice but lack LAPI’s level of detail and actionable guidance. For instance, Chain-of-Thought misses lifestyle advice, and RAG-Based Few-Shot doesn’t emphasize continuous patient-doctor communication as strongly as LAPI.
6 Conclusion
In this work, we introduce a novel approach for designing LLM-based professional service agents tailored for medical Q&A services. Our method combines theory-guided task planning and pragmatic entropy to ensure professional-aligned communication for ethical, patient-centric responses. Experiments show that our method outperforms baseline techniques in fluency, empathy, patient-centricity, and ROUGE-L scores. This research highlights the potential of enhancing AI chatbot communication, particularly in healthcare, and encourages further development of more effective methods in future work.